OpenStudy (anonymous):

Evaluate 48 ÷ (6 + 2) • 3 4 18 14 30

OpenStudy (dan815):

u know of BedMAS?

OpenStudy (dan815):

brackets exponents division and multiplication addition and subtraction

OpenStudy (anonymous):

No I Was Taught PEMDAS

OpenStudy (dan815):

lol ok wow..

OpenStudy (dan815):

so do it in that order

OpenStudy (anonymous):

parenthesis exponents multiplication division addition subtraction

OpenStudy (dan815):

yes yes samething ok

OpenStudy (dan815):

now do it

OpenStudy (anonymous):

6+2=8 48 ÷ 8• 3 28 ÷ 48

OpenStudy (dan815):

hmm something went crazy on the last step

OpenStudy (anonymous):

is tht right cause i get an answer that isn;t in the answer choices.

OpenStudy (dan815):

nope

OpenStudy (anonymous):

24

OpenStudy (anonymous):

how about this So 48÷(6+2)• 3 =48÷(8)• 3 =6• 3 =18

OpenStudy (dan815):

yes that is right

OpenStudy (anonymous):

thnxs

OpenStudy (dan815):

thanks